### v1

# Soft Sequence Policy Optimization: Bridging GMPO and SAPO

Abstract
A significant portion of recent research on Large Language Model (LLM) alignment focuses on developing new policy optimization methods based on Group Relative Policy Optimization (GRPO). Two prominent directions have emerged: (i) a shift toward sequence-level importance sampling weights that better align with the sequence-level rewards used in many tasks, and (ii) alternatives to PPO-style clipping that aim to avoid the associated loss of training signal and entropy collapse. Recent work, such as Soft Adaptive Policy Optimization (SAPO), reformulates the Scopic objective within the GRPO framework and achieves both sequence coherence and token adaptivity. Geometric-Mean Policy Optimization (GMPO) leverages token-wise ratio clipping within sequence importance sampling weights. Building on these ideas, this work proposes a new objective that promotes effective policy exploration while maintaining training stability. Specifically, we introduce Soft Sequence Policy Optimization, an off-policy reinforcement learning objective that incorporates soft gating functions over token-level probability ratios within sequence-level importance weights.
1 Introduction
Reinforcement learning (RL) has become a central ingredient in enhancing the reasoning abilities of large language models (LLMs), particularly for tasks requiring long chains of thought (CoT), multi-step decision making, and delayed rewards. While supervised fine-tuning excels at imitating local patterns present in static datasets, RL enables optimization over entire generated sequences, making it especially attractive for complex reasoning, program synthesis, and decision-centric language modeling.
Among modern RL approaches for LLM alignment, group-based policy optimization methods have emerged as a practical and effective recipe. These methods sample multiple candidate responses per input prompt and normalize sequence-level rewards within each group, yielding a reliable prompt value estimate without relying on an auxiliary critic network. Representative algorithms such as Group Relative Policy Optimization (GRPO) [Shao et al., 2024] and REINFORCE Leave-One-Out (RLOO) [Ahmadian et al., 2024] demonstrate that relative comparisons within a group can significantly stabilize training, reduce computational overhead, and improve reasoning performance across a wide range of domains.
Despite their empirical success, existing group-based methods face important limitations when deployed at scale. In particular, off-policy learning becomes unavoidable in realistic training pipelines. As model sizes increase, architectures incorporate sparsity (e.g., Mixture-of-Experts), and generated sequences become longer, large rollout batches are required to fully utilize modern hardware. To improve sample efficiency, these batches are typically partitioned into multiple mini-batches for gradient updates. This practice inevitably induces an off-policy setting, where responses are sampled from a behavior policy while updates are applied to a newer policy, a regime in which GRPO-like algorithms increasingly dominate current practice.
In off-policy group-based optimization, policy updates are usually weighted by importance sampling (IS) ratios between the current and behavior policies. These ratios are well known to suffer from high variance, especially for long sequences where token-level likelihood ratios compound multiplicatively. Uncontrolled IS weights can destabilize training and ultimately degrade performance. Two broad strategies have emerged to mitigate this issue. First, many methods apply hard clipping to large importance weights, limiting their influence on the gradient [Schulman et al., 2017, Shao et al., 2024, MiniMax et al., 2025]. While effective at reducing variance, hard clipping introduces a difficult tradeoff: aggressive clipping improves stability but reduces sample efficiency and limits exploration [Chen et al., 2022, Dwyer et al., 2025], whereas loose clipping preserves learning signal at the cost of noisy and brittle updates.
Second, in accordance with insights from RLOO and motivated by the prevalence of sequence-level rewards in LLM alignment (e.g., RLHF and RLVR), recent work proposes enforcing sequence-level coherence of importance weights rather than treating token-level contributions independently. This perspective motivates objectives such as Group Sequence Policy Optimization (GSPO) [Zheng et al., 2025] and Geometric-Mean Policy Optimization (GMPO) [Zhao et al., 2025], which replace arithmetic aggregation of token-level ratios with multiplicative or geometric formulations that better respect the structure of sequence probabilities. A key innovation of GMPO is the combination of sequence-level weighting with token-level probability ratio clipping.
However, existing solutions remain incomplete. Sequence-coherent objectives, such as GSPO, improve stability but do not fully address the interaction between off-policy learning and entropy-regularized objectives commonly used in modern RL for LLMs. In parallel, soft policy optimization methods — such as Soft Adaptive Policy Optimization (SAPO) [Gao et al., 2025] — highlight the importance of entropy-aware objectives that smoothly interpolate between exploitation and exploration, but they are not explicitly designed for sequence-level optimization.
In this paper, we propose Soft Sequence Policy Optimization (SSPO), a new off-policy reinforcement learning objective that unifies and extends insights from sequence-level and soft policy optimization. SSPO introduces a soft, sequence-level weighting mechanism with token-level weight attenuation that controls importance sampling variance without resorting to hard clipping, while preserving coherent credit assignment to entire responses. By operating at the sequence level and incorporating entropy-aware regularization, SSPO achieves a more favorable bias–variance tradeoff than prior approaches in off-policy group-based RL.
The main contributions of this work are as follows:
1.
We propose Soft Sequence Policy Optimization (SSPO), a sequence-level off-policy objective that combines geometric aggregation with smooth token-level gating.
2.
We provide theoretical analysis of the resulting objective, highlighting its bias–variance properties and its relationship to GRPO, GSPO, GMPO, and SAPO.
3.
We outline the empirical evaluation protocol for SSPO on RL-based LLM alignment tasks to analyze its stability and sample efficiency relative to existing baselines.
2 Related Works
Group-based policy optimization has become a strong baseline for reinforcement learning with large language models, particularly in alignment and reasoning tasks. Methods such as GRPO [Shao et al., 2024] and RLOO [Ahmadian et al., 2024] estimate advantages by sampling multiple responses per prompt and computing relative rewards within each group, avoiding an auxiliary critic network. GRPO further enables off-policy updates leveraging importance-weighted updates, making it a practical default in large-scale RL pipelines and inducing various modifications [Yu et al., 2025, Liu et al., 2025b, a].
In realistic training setups, off-policy learning is unavoidable due to large rollout buffers and mini-batch gradient steps. However, applying token-level importance sampling ratios to sequence-level rewards introduces a policy–reward mismatch that can destabilize training. Recent work addresses this issue by enforcing sequence-level coherence in policy updates. GSPO [Zheng et al., 2025] and GMPO [Zhao et al., 2025] substitute arithmetic mean of token-level ratios with sequence-consistent formulations, improving stability and variance control.
Hard clipping of importance ratios, originally introduced in PPO [Schulman et al., 2017], remains a common strategy for variance reduction in off-policy optimization [Shao et al., 2024, MiniMax et al., 2025]. While effective, clipping induces a fundamental bias–variance tradeoff: aggressive clipping improves stability but harms sample efficiency and exploration, whereas loose clipping leads to noisy updates [Chen et al., 2022, Dwyer et al., 2025]. Several works explore alternatives, including asymmetric clipping [Yu et al., 2025] and gradient-preserving approaches [Su et al., 2025a], but these methods remain sensitive to hyperparameters.
An alternative line of work advocates for soft, entropy-aware objectives that avoid hard clipping. SAPO [Gao et al., 2025] and related soft trust-region methods [Dwyer et al., 2025] demonstrate improved robustness in off-policy settings by smoothly attenuating large importance ratios. Recent work also explores global soft constraints on policy distributions, such as Entropy Ratio Clipping [Su et al., 2025b], which uses the ratio of policy entropies between updates to constrain distributional drift beyond standard importance clipping. However, these approaches are primarily developed for token-level objectives and do not explicitly address sequence-level credit assignment in group-based learning.
3 Preliminaries
Since we do not modify the additive KL-divergence regularization term in the loss, we omit it for brevity.
Notation
Following standard conventions in the literature, an autoregressive language model parameterized by is defined by a policy . denotes a query and denotes the set of queries. Given a response to a query , the likelihood of under the policy is defined as
where denotes the number of tokens in . Each query–response pair is assigned a reward by a verifier .
Group Relative Policy Optimization (GRPO)
Building on Proximal Policy Optimization (PPO) [Schulman et al., 2017], Group Relative Policy Optimization (GRPO) [Shao et al., 2024] enables off-policy learning by leveraging samples generated by an older (behavior) policy . Similar to PPO, GRPO employs a clipping mechanism to constrain policy updates within a proximal region around the behavior policy. The key difference is that GRPO replaces computationally and memory-intensive advantage estimators based on a critic network —- whose reliability may be questionable — with a relative advantage computed among responses to the same query.
Here, denotes the group size, i.e., the number of responses generated for a single query. The importance ratio and the advantage for token are defined as
The notation is valid since all tokens in the sequence share the same advantage.
Group Sequence Policy Optimization (GSPO)
Group Sequence Policy Optimization (GSPO) [Zheng et al., 2025] identifies a primary problem of GRPO: a mismatch between the unit of optimization and the unit of reward. In GRPO, importance sampling weights for off-policy correction and clipping are applied at the token level, whereas the advantage remains constant across all tokens in a sequence. In contrast, GSPO performs optimization — and in particular clipping — directly at the sequence level. Its objective is given by
where the importance ratio is computed based on the sequence likelihood, with length normalization to control variance [Zheng et al., 2023]:
GSPO has been shown to improve training stability for large models and to increase sample efficiency, even as the fraction of clipped tokens grows.
Geometric-Mean Policy Optimization (GMPO)
Geometric-Mean Policy Optimization (GMPO) [Zhao et al., 2025] combines response-level importance sampling with token-level probability ratio clipping. GMPO adopts the same sequence-length normalization as GSPO but is motivated by a different intuition: reducing sensitivity to outliers and variance in token-level importance sampling weights by replacing the arithmetic mean over tokens in the GRPO objective with a geometric mean. Specifically, GMPO optimizes the following objective:
This geometric aggregation improves robustness to tokens with disproportionately large importance weights and reduces the range of the loss values, leading to more stable updates. The authors further observe that token-level clipping is less aggressive than sequence-level clipping and more effectively constrains the range of importance sampling ratios.
Soft Adaptive Policy Optimization (SAPO)
In existing group-based policy optimization methods, high variance in importance sampling weights is typically mitigated through hard clipping. This introduces a trade-off between update stability and sample efficiency and may also reduce exploration abilities. As an alternative, Soft Adaptive Policy Optimization (SAPO) [Gao et al., 2025] employs a smooth, temperature-controlled gating mechanism that preserves the learning signal for all sampled tokens. The SAPO objective is defined as:
where
Here, and denote the temperatures for positive and negative advantages, respectively, and is the sigmoid function.
By construction, the sigmoid-shaped gating function preserves gradients during on-policy updates while remaining bounded. Attenuating the contribution of outlier tokens, rather than truncating it entirely, allows SAPO to maintain informative learning signals throughout training and induces a continuous trust region.
4 Soft Sequence Policy Optimization
While Soft Adaptive Policy Optimization (SAPO) preserves token-level adaptivity through smooth gating, it is sequence-coherent only under two mild assumptions: (i) small policy updates, i.e., , and (ii) low intra-sequence dispersion of importance ratios, . These assumptions may be violated in practice, particularly for long sequences or during the early stages of training. Our goal is therefore to redesign the SAPO objective to relax these assumptions while retaining token-level adaptivity.
A key challenge is balancing stability and expressiveness. Following SAPO, we retain distinct temperatures for positive and negative advantages, with . This choice is motivated by two complementary observations. First, DAPO [Yu et al., 2025] shows that increasing the effective upper bound on policy updates (Clip-Higher) improves exploration and mitigates entropy collapse. Second, SAPO observes that negative-advantage tokens tend to introduce greater instability, as their gradients increase the probability mass of many unsampled and often irrelevant tokens. As noted in prior work, positive advantages increase the sampled token’s logit while suppressing unsampled logits, whereas negative advantages have the opposite effect.
In the absence of clipping, the geometric mean of token-wise importance ratios recovers the length-normalized sequence-level importance ratio used in GSPO. GMPO can thus be interpreted as a relaxation of sequence-level policy optimization that retains token-level adaptivity.
To better align token-level adaptivity with sequence-level coherence, we propose Soft Sequence Policy Optimization (SSPO). SSPO aggregates token-level gating functions using a geometric mean, yielding the following objective:
Here, is the sigmoid-based gating function defined in Eq. (7). The absolute value used in GMPO is omitted since implies for any positive temperature.
5 Experiments
We evaluate SSPO on representative RL-based LLM alignment tasks requiring complex reasoning. The evaluation protocol is designed to assess training stability, variance of importance weights, and sample efficiency under off-policy updates. We benchmark SSPO against standard GRPO, GSPO, GMPO, and SAPO using mathematical reasoning datasets. The primary metrics include task performance and empirical measurements of entropy collapse.
6 Conclusion
This paper introduced Soft Sequence Policy Optimization (SSPO), a sequence-level off-policy objective for reinforcement learning with large language models. The proposed formulation combines geometric aggregation of token-level contributions with smooth sigmoid-based gating, aligning the unit of reward with the unit of importance correction while avoiding hard clipping.
By attenuating extreme importance ratios instead of truncating them, SSPO aims to achieve a more favorable bias–variance tradeoff and improved training stability in group-based learning settings. The framework provides a unified perspective connecting GRPO, GSPO, GMPO, and SAPO, and establishes a foundation for further empirical investigation of sequence-coherent soft policy optimization methods.
Computational Resources
The research was carried out using the MSU-270 supercomputer of Lomonosov Moscow State University.
Acknowledgements
This work was supported by the Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier 000000C313925P4H0002; grant No 139-15-2025-012).